# Thermal Casimir effect in $\kappa$-Minkowski space-time

### v1

@@ 1 Introduction
Non-commutative (NC) space-times inherently introduce a minimal length scale, making it of intrinsic interest to study the implications of this scale in physical phenomena. One such phenomenon where the length scale plays a significant role is the Casimir effect [3, 4]. It was shown that when two uncharged conducting plates are placed parallel in a small separation, they experience an attractive force [3, 4, 5, 6, 7]. Modifications to the Casimir effect due to different dimensions of space-time, the presence of medium between plates, the thermal effect, etc., using real scalar field theory have been studied extensively [5, 4, 6, 8]. High-precision experiments have quantified the Casimir force between parallel plates at micrometer-scale separations [9, 10, 11]. Rigorous theoretical and experimental studies have also explored the Casimir effect, beyond parallel conducting plates, in various geometric configurations [12, 13, 14].
@@ 3 Deformed finite temperature Casimir effect
Now, consider two parallel plates fixed at $z=0$ and $z=L$ and by imposing the Dirichlet boundary condition, we find the momentum component along the $z$ direction to be $k_{m}=\frac{m\pi}{L}$ where $m=1,2,$.... Thus, we get $k=\sqrt{k_{\perp}^{2}+k_{m}^{2}}$ and the $dk_{m}$ becomes $\frac{dk_{m}}{2\pi}\to\frac{1}{L}\displaystyle\sum_{m=1}^{\infty}$. By rewriting the volume $V$ in terms of $L$ and area $A$, we get the Casimir free energy per unit area, i.e., $\mathcal{F}=F/A$, as
@@
First, we evaluate the commutative term by expanding the logarithmic term using $\ln{(1-x)}=-\displaystyle\sum_{n=1}^{\infty}\frac{x^{n}}{n}$ and we get $\mathcal{F}_{T}^{~(0)}=-\frac{1}{2\pi\beta^{3}}\displaystyle\sum_{m=1}^{\infty}\displaystyle\sum_{n=1}^{\infty}\frac{1}{n}\displaystyle\int_{m\pi/L}^{\infty}d\bar{k}~\bar{k}e^{-n\beta\bar{k}}$, where $\bar{k}=\sqrt{k_{\perp}^{2}+m^{2}\pi^{2}/L^{2}}$. Carrying out this integral and evaluating the summation over $n$, we arrive at the standard expression for $\mathcal{F}_{T}^{(0)}$ as [33]
@@ Acknowledgements
Figure 6: Integration contour $C$ containing the poles at $z=i\omega n$ (Matsubara frequencies)

### v2

@@ 1 Introduction
Non-commutative (NC) space-times inherently introduce a minimal length scale, making it of intrinsic interest to study the implications of this scale in physical phenomena. One such phenomenon where the length scale plays a significant role is the Casimir effect [3, 4]. It was shown that when two uncharged conducting plates are placed parallel in a small separation, they experience an attractive force [3, 4, 5, 6, 7]. Modifications to the Casimir effect due to different dimensions of space-time, the presence of medium between plates, the thermal effect, etc., using real scalar field theory have been studied extensively [5, 4, 6, 8]. High-precision experiments have quantified the Casimir force between parallel plates at micrometer-scale separations [9, 10, 11]. Rigorous theoretical and experimental studies have also explored the Casimir effect, beyond parallel conducting plates, in various geometric configurations [12, 13, 14]. Casimir effect for various other configurations, such as piston configuration with imposed curved boundary conditions and in the presence of compact dimensions, has also been extensively studied [15].
@@ 3 Deformed finite temperature Casimir effect
Now, consider two parallel plates fixed at $z=0$ and $z=L$ and by imposing the Dirichlet boundary condition, we find the momentum component along the $z$ direction to be $k_{m}=\frac{m\pi}{L}$ where $m=1,2,$.... Thus, we get $k=\sqrt{k_{\perp}^{2}+k_{m}^{2}}$ and $dk_{m}$ becomes $\frac{dk_{m}}{2\pi}\to\frac{1}{L}\displaystyle\sum_{m=1}^{\infty}$. By rewriting the volume $V$ in terms of $L$ and area $A$, we get the Casimir free energy per unit area, i.e., $\mathcal{F}=F/A$, as
@@
Note that the explicit form of $k$ changes as the imposed boundary condition changes (as shown in [15]). Similarly, if we consider a compact dimension (say along the $y$ direction), then we obtain $k=\sqrt{k_{x}^{2}+\left(\frac{2\tilde{m}\pi}{R}\right)^{2}+\frac{m^{2}\pi^{2}}{L^{2}}}$, where $\tilde{m}$ is an integer and $R$ is the length of the compact dimension. All these modifications seen in $k$ due to changes in the boundary conditions or due to the presence of compact dimensions, lead to the modification of $\mathcal{F}_{T}$, given in Eq.(3.9). Hence, changes in the boundary conditions or the presence of compact dimensions affect the final expression for the thermal Casimir energy, even under $\kappa$-deformation. Such modifications can also be observed if we consider non-trivial field configurations which induce anti-periodic boundary conditions on the bosonic field, as in [38].
First, we evaluate the commutative term of Eq.(3.9) by expanding the logarithmic term using $\ln{(1-x)}=-\displaystyle\sum_{n=1}^{\infty}\frac{x^{n}}{n}$ and we get $\mathcal{F}_{T}^{~(0)}=-\frac{1}{2\pi\beta^{3}}\displaystyle\sum_{m=1}^{\infty}\displaystyle\sum_{n=1}^{\infty}\frac{1}{n}\displaystyle\int_{m\pi/L}^{\infty}d\bar{k}~\bar{k}e^{-n\beta\bar{k}}$, where $\bar{k}=\sqrt{k_{\perp}^{2}+m^{2}\pi^{2}/L^{2}}$. Carrying out this integral and evaluating the summation over $n$, we arrive at the standard expression for $\mathcal{F}_{T}^{(0)}$ as [34]
@@ 5 Conclusion
However, the calculations become slightly more intricate if we consider a fermionic field, where we need to use the MIT bag model to set boundary conditions, as in [46], to obtain the corresponding eigen frequency for calculating the Casimir energy between parallel plates. Since the fermionic field satisfies the anti-periodic boundary conditions, the resulting Matsubara frequency modes are given by $\omega_{n}=\frac{(2n+1)\pi}{\beta}$ [36, 37]. Additionally, we need to consider the correction terms coming from the $\kappa$-deformed Lagrangian for Dirac field [47]. Incorporating these modifications in our study changes the final results of thermal Casimir energy, Casimir force, and entropy. Studies along these directions are in progress and will be reported soon.
@@ Acknowledgements
Figure 6: Integration contour $C$ containing the poles at $z=i\omega_{n}$ (Matsubara frequencies)